OpenStudy (anonymous):

Relationship B has a lesser rate than Relationship A. The graph represents Relationship A. Which table could represent Relationship B? A. Time (weeks) 3 6 8 10 Plant growth (in.) 2.25 4.5 6 7.5 B. Time (weeks) 3 6 8 10 Plant growth (in.) 4.8 9.6 12.8 16 C. Time (weeks) 3 4 6 9 Plant growth (in.) 5.4 7.2 10.8 16.2 D. Time (weeks) 3 4 6 9 Plant growth (in.) 6.3 8.4 12.6 18.9

OpenStudy (anonymous):

@lovielove456767890

OpenStudy (anonymous):

plz help me

OpenStudy (danjs):

Need the graph, to get the slope of the graph

OpenStudy (danjs):

then compare it to the rate of change in eachof the tables

OpenStudy (anonymous):

sorry forgot the graph

OpenStudy (danjs):

slope of graph = (6-3)/(4-2) = 3/2 Rate of change of Graph is 3/2

OpenStudy (anonymous):

@dtan5457 plz help

OpenStudy (danjs):

Table A (4.5-2.25)/(6-3) = 0.75 rate of change

OpenStudy (anonymous):

YOU WERE SUPPOSE TO MESSAGE ME @TeenWolfGirl

OpenStudy (anonymous):

I did

OpenStudy (danjs):

looks like table A is the one

OpenStudy (anonymous):

mad or nah

OpenStudy (danjs):

Graph = 1/2 Table A = 0.75

OpenStudy (danjs):

Table B (9.6-4.8)/6-3) = 1.6

OpenStudy (danjs):

Table C = 1.8

OpenStudy (danjs):

Table A has a smaller rate of change than the Graph